• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 22:33:18 on February 25, 2011

    Entry number 348673


    keyword=DAQ Deadtime Issue

    After leaving the hall, Tai noticed that the DAQ deadtime was high for
    cosmics.  We were taking only 14 Hz of T1 and 15 Hz of T3 with 1.5 Hz of 
    T6.  The majority of these are EDTM signals, which should cause no real 
    deadtime.  The online deadtime monitor indicates that the deadtime is 
    about 35%, but the end of run indicates about 62% or so.
    I changed CODA configs to what we ran for D-threshold and the deadtime
    persisted indicating the problem is probably with one or both HRS's.  I
    then took some T8 runs with ps8 = 4 and 1.  The online DT monitor
    indicated that the DT was 0%, but the end of run file indicated that it 
    was about 36% and 50% respectively.  Previously the two were fairly close
    to each other.
    I also took T1 and T3 only runs.  The online monitor again indicated 
    no deadtime, but the end-of-run showed about 40%.
    We thought perhaps there was some extra counting due to the EDTM for
    left and right arm coming near each other since we had about 9 Hz of 
    EDTM coincidences.  I changed  the EDTM delays to remove the coincidence
    but the deadtime persisted.
    David is in the hall checking the HRS's for missing gates.

    A copy of this log entry has been emailed to: lselvy,danez,yez,rom,kalyan